X + 2y - z = 0
X + y - 2z = 3
2x - z = 0

Answer by mananth(16946)

X + 2y - z = 0
X + y - 2z = 3
2x - z = 0


2x -z=0
2x=z
X + y - 2z = 3
X + y - 2*2x = 3
y-3x = 3......................4

X + 2y - z = 0
X + 2y - 2x = 0
2y-x=0.........................5

solve 4 &5 to get y=-3/5
plug y and get x &z